A company has determined that the price and monthly demand of one of its products are related by the equation D = (4000 - p)2/3 where p is the price per unit in dollars and D is the monthly demand in units. The associated fixed costs are $850 per month, and the variable costs are $3,687.5 per unit. What is the optimal number of units (rounded to the next whole number) that should be produced and sold each month?
Answer #1

Q=(4000-P)2/3

P=4000-Q3/2

TR=4000Q-Q5/2

profit is maximised when MR=MC and we know MC=change in variable cost=3687.5

marginal revenue=dTR/dQ=4000-5/2Q3/2=3687.5

5/2Q3/2=312.5

Q3/2=312.5*2/5=125

Q=25 Should be produced and sold out